# Precise calculation of the EFT likelihood with primordial non-Gaussianities

### v1

@@ 4 Precise calculation of the EFT likelihood
Starting from this stage, we will formally proceed with the detailed calculation of the conditional likelihood under non-Gaussian initial conditions. As seen from eqs. (3.11) and (3.13), the inclusion of PNG introduces the same interaction term to both partition functions. In [23], the authors considered the case where only the noise in the galaxy power spectrum $P_{\epsilon_{g}}$ is present and argued that, in this context, the non-Gaussian term contributes identically to the extra terms in both likelihoods. Consequently, when evaluating the ratio of the two — corresponding to the condition likelihood — their contributions precisely cancel, yielding the same result as was previously obtained. Thus, in order to obtain nontrivial results, in this paper, we will focus on the case where both $P_{\epsilon_{g}}$ and $P_{\epsilon_{g}\epsilon_{m}}$ are present. We will illustrate through detailed calculations that the additional contributions to the two likelihoods originating from the saddle points do not cancel each other completely. Moreover, when accounting for the contributions around the saddle points, a field dependent, irreducible factor will also be introduced in the final conditional likelihood.
@@ 4.1 Set up the likelihood expressions
We begin by concerned with the explicit formulae of the matter likelihood and joint likelihood. As discussed in [24], following an appropriate reformulation, the expression for the likelihood can be written as a path integral in three-dimensional Euclidean space. Firstly, the matter likelihood eq. (3.14) can be reformulated as follows
@@
$\mathcal{T}^{abc}=\left\{\begin{array}[]{ll}(2\pi)^{3}\delta_{D}(\boldsymbol{k}+\boldsymbol{p}_{12})f_{NL}B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\;,&\text{for}\;a=b=c=2\,,\\ 0\;,&\text{otherwise }\,,\end{array}\right.$ (4.5)
@@
In this context, we have preformed a similar package of the integral variables $\boldsymbol{\phi}_{g}=(X_{g},X,\delta_{\rm in})^{T}$. The action at this stage can be expressed as
@@
The expressions for the matrices can be read off in the same manner
@@
$\mathcal{T}_{g}^{abc}=\left\{\begin{array}[]{ll}(2\pi)^{3}\delta_{D}(\boldsymbol{k}+\boldsymbol{p}_{12})f_{NL}B(\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2})\;,&\text{for}\;a=b=c=3\,,\\ 0\;,&\text{otherwise }\,.\end{array}\right.$ (4.9)
The primary objective of this paper is to compute the two likelihoods eq. (4.1), eq. (4.6), as well as the conditional likelihood defined by their ratio. Evidently, this requires us to employ the saddle-point expansion method discussed in section 2 to perform the calculation, and we will demonstrate that, the implementation of PNG leads to nontrivial additional contributions to the conditional likelihood. In accordance with the previous spirit, we should decompose the path integral into contributions from the saddle points and their vicinity. We will now proceed to discuss these contributions individually.
@@ 4.2 The contributions from the saddle points
It is evident that both of these saddle-point equations are quadratic matrix equations, and therefore their exact solutions typically involve the application of some eigenvalue analysis techniques. Fortunately, in our case, the matrices $\mathcal{T}$ and $\mathcal{T}_{g}$ can be treated as perturbations, as they are associated with PNG and are therefore proportional to $f_{\rm NL}$. This allows us to obtain perturbative solutions of these equations. And this is precisely what we anticipated, since our focus is on the subtle corrections to the likelihood terms that emerge upon implementing PNG. We will then discuss separately the computation of the matter likelihood and the joint likelihood, along with their leading-order contributions corresponding to PNG.
@@
By iterating employing this method, we can obtain solutions to arbitrary higher orders. It is straightforward to observe that, after incorporating the leading-order PNG, the corrected saddle-point solution compared to eq. (4.13) effectively corresponds to a shift of the matter current $X_{\rm cl}\rightarrow X_{\rm cl}+\frac{iBf_{\rm NL}\delta^{2}}{D_{1}^{3}}$, and this correction is hierarchically suppressed by powers of $f_{\rm NL}$. This property enables us to carry over the discussion and results on the saddle-point contribution from [23], and to explore additional terms induced by the shift of $X$. Specifically, the terms in the action that give rise to corrections at order $f_{\rm NL}$ due to this adjustment are
@@
The aforementioned terms correspond to the supplementary contributions to the matter likelihood upon the introduction of PNG. Note that in calculating these contributions, we must not only consider the terms in the action that depend on $X$, but also take into account the non-Gaussian effective interaction. By employing the same method, we can also extend this calculation to arbitrary higher orders.
@@
We now examine a more complicated scenario concerning the joint likelihood. Given the complexity of its formulation, we will first consider the simplest case, where $P_{\epsilon_{g}\epsilon_{m}}=0$, and subsequently restore the general case. By assuming the same field-decomposition ansatz and following an analgous derivation process, we can show that
$\bar{\boldsymbol{\phi}}_{g}=\boldsymbol{\phi}_{g,0}+\boldsymbol{\phi}_{g,1}\,,\;\boldsymbol{\phi}_{g,1}=-\frac{1}{2}(\mathcal{M}_{g}^{-1})^{ab}\mathcal{T}_{g}^{bcd}\boldsymbol{\phi}_{0,g}^{c}\boldsymbol{\phi}_{0,g}^{d}=\begin{pmatrix}0\\ \frac{iBf_{\rm NL}\delta^{2}}{2D_{1}^{3}}\\ 0\end{pmatrix}\,,$ (4.20)
@@
By comparing eqs. (4.20) and (4.21), we can observe that after recovering $P_{\epsilon_{g}\epsilon_{m}}$, in addition to reproducing the same shift as before, all remaining terms are at least proportional to $P_{\epsilon_{g}\epsilon_{m}}$ or $f_{\rm NL}$. On the other hand, from our previous discussion, we know that $P_{\epsilon_{g}\epsilon_{m}}\sim k^{2}$, implying that its contribution is suppressed at large scales. This enables us to treat it as a perturbative parameter similar to $f_{\rm NL}$. Note that the leading-order saddle-point solution [24]
@@
itself inherently involves $P_{\epsilon_{g}\epsilon_{m}}$. Therefore, if we include the $P_{\epsilon_{g}\epsilon_{m}}$-related components in $\boldsymbol{\phi}_{g,1}$, many of the resulting terms naturally correspond to higher-order contributions. In the present work, our objective is to isolate the leading-order contributions proportional to $P_{\epsilon_{g}\epsilon_{m}}$ and $f_{\rm NL}$. Accordingly, it is reasonable to approximate the first-order perturbed solution as follows
@@
$X_{g,\rm cl}$ $\rightarrow X_{g,\rm cl}-\frac{iBf_{\rm NL}P_{\epsilon_{g}\epsilon_{m}}\delta^{2}}{2P_{\epsilon_{g}}D_{1}^{3}}\,,$ (4.25)
$\,X_{\rm cl}$ $\rightarrow X_{\rm cl}+\frac{Bf_{\rm NL}}{2D_{1}^{3}}\left(i\delta^{2}-2P_{\epsilon_{g}\epsilon_{m}}\delta X_{g,\rm cl}-\frac{ib_{1}P_{\epsilon_{g}\epsilon_{m}}\delta^{2}}{P_{\epsilon_{g}}}\right)\,,$ (4.26)
@@
×$\times\exp\left\{-\int_{\boldsymbol{k}}P_{\epsilon_{g}\epsilon_{m}}X_{g}X+\frac{f_{\rm NL}}{3!}\int_{\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2}}B(\delta_{\rm in})^{3}\right\}\bigg{|}_{\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}}\,,$ (4.27)
@@
Moreover, all other terms are at least at order $\mathcal{O}(f_{\rm NL}^{2})$ or $\mathcal{O}(P_{\epsilon_{g}\epsilon_{m}}^{2})$. At the same time, one might be concerned that certain terms in the action contain explicit factors of the imaginary unit $i$, potentially rendering the expression complex. However, this is not the case: due to the presence of $i$ in the definition of $X_{g,\rm cl}$ itself in eq. (4.22), all imaginary components in $S_{g}[\bar{\phi}_{g}]$ cancel out exactly, ensuring that the resulting expression is manifestly real — consistent with the physical interpretation of probability.
A brief comparison of the expressions for the two likelihoods with those under Gaussian initial conditions reveals certain underlying structures. We can observe that, when $P_{\epsilon_{g}\epsilon_{m}}=0$, the contributions of PNG to both likelihoods are exactly the same, consistent with the discussion presented in [23]. Conversely, when the cross stochastic term is introduced, its contributions to the joint likelihood prevent this cancellation when computing the ratio of the two likelihoods, thereby imparting non-vanishing contributions to the conditional likelihood. The relationship between these noise terms and the likelihood expressions is of considerable interest, we will discuss it in detail in the following subsection.
@@ 4.3 The contributions around the saddle points
We now turn to the section on precise calculations — specifically, employing the saddle-point expansion method to compute the contributions around the saddle points. To achieve this, we decompose the fields $\boldsymbol{\phi}$ and $\boldsymbol{\phi}_{g}$ into their saddle-point solutions and some small fluctuations around them
@@
It is evident that if the two eigenvalues are real, they must also be positive, as the growth factor $D_{1}$ can be consistently expected to be positive. Therefore, there are no concerns regarding negative-mode solutions at this stage, so we can safely proceed with the saddle-point expansion method and apply the Gaussian integration formula, leading to the following result
@@
Here, when plugging $S[\bar{\boldsymbol{\phi}}]$ we have already incorporated the additional contributions arising from the inclusion of PNG. And in this case, it is evident that although the eigenvalues $m_{n}$ exhibit explicit dependence on the matter field $\delta$, such dependence cancels out in their product. Consequently, the resulting matter likelihood precisely matches that obtained in [24]. As established in the previous discussions, this field-independent prefactor will be consistently absorbed into the normalization factor associated with the definition of functional Dirac delta function, and therefore it does not yield any nontrivial contribution to the matter likelihood.
@@
Even though we have employed such a redefinition to simplify our calculation, $S_{g}^{\prime\prime}[\bar{\boldsymbol{\phi}}_{g}]$ remains a three-dimensional matrix and may still process negative or complex eigenvalues, thereby rendering the corresponding functional integral ill-defined. Accordingly, when computing the joint likelihood, it is essential to follow the strategy employed in section 2, complexifying the action and identifying the combination of the steepest descent contours in the complex plane that is homologous to the original integration contour (the real line in the field space). The full path integral can then be approximated by summing over all the contributions from these complex contours, which ensures a well-defined and convergent result. All relevant technical details have been provided in the preceding sections, and here we directly employ the gradient flow equation to determine the corresponding steepest descent contours. Assuming that the same steps as previously outlined have already been performed, the gradient flow equation in this case can be written as
@@
where we have used the fact that $\bar{\boldsymbol{\phi}}_{g}$ is independent of $u$, since $u$ simply serves as a parameter labeling different integration contours through the saddle point in the complex plane. At this time, the gradient flow equation is modified from its previous form eq. (2.7) by the addition of a non-Gaussian term, which is precisely equivalent to implementing the variable shift eq. (4.42) (note that $P_{\rm in}^{\prime\prime-1}$ is real). Therefore, by implementing this shift, one can construct the same eigenequation and convert it into a Hermitian eigenequation as in section 2, with the expressions remaining identical to those derived previously. Consequently, the sum over these positive and real eigenvalues $m_{n}$ can be obtained
@@
From this point forward, we will use the simplified notation $P_{\rm in}^{\prime\prime-1}$, with the understanding of its dependence on the fields $\delta$ and $\delta_{g}$. With these elements established, we are now ready to compute the corresponding joint likelihood
@@
We can see that, the resulting expression contains not only the saddle-point contributions but also an additional field-dependent prefactor. At this stage, the prefactor exhibits an explicit dependence on both the matter and galaxy fields, and cannot be eliminated by setting $P_{\epsilon_{g}\epsilon_{m}}=0$. In other words, once the contributions of PNG are taken into account, the precise calculation process introduces a field-dependent term that cannot be absorbed into the normalization procedure. Consequently, this term appears explicitly in the final expression for the conditional likelihood. Its potential implications will examined in detail in the following subsection.

### v3

@@ 4 Precision calculation of the EFT likelihood
Starting from this stage, we will formally proceed with the detailed calculation of the conditional likelihood under non-Gaussian initial conditions. As seen from eqs. (3.11) and (3.13), the inclusion of PNG introduces the same interaction term to both partition functions. In [23], the authors considered the case where only the noise in the galaxy power spectrum $P_{\epsilon_{g}}$ is present and argued that, in this context, the non-Gaussian term contributes identically to the extra terms in both likelihoods. Consequently, when evaluating the ratio of the two, corresponding to the conditional likelihood, their contributions precisely cancel, yielding the same result as was previously obtained. Thus, in order to obtain nontrivial results, in this paper, we will focus on the case where both $P_{\epsilon_{g}}$ and $P_{\epsilon_{g}\epsilon_{m}}$ are present. We will illustrate through detailed calculations that the additional contributions to the two likelihoods originating from the saddle points do not cancel each other completely. Moreover, when accounting for the contributions around the saddle points, a field dependent, irreducible factor will also be introduced in the final conditional likelihood.
@@ 4.1 Set up the likelihood expressions
We begin by focusing on the explicit formulae of the matter likelihood and joint likelihood. As discussed in [24], following an appropriate reformulation, the expression for the likelihood can be rewritten as a path integral in three-dimensional Euclidean space. Firstly, the matter likelihood eq. (3.14) can be reformulated as follows
@@
[…]
@@
In this context, we have preformed a similar packaging of the integral variables $\boldsymbol{\phi}_{g}=(X_{g},X,\delta_{\rm in})^{T}$. The action at this stage can be expressed as
@@
The expressions for the matrices can be derived in the same manner
@@
[…]
The primary objective of this paper is to compute the two likelihoods eq. (4.1), eq. (4.6), as well as the conditional likelihood defined by their ratio. Evidently, this requires us to employ the saddle-point expansion method discussed in section 2 to perform the calculation. We will demonstrate that the implementation of PNG leads to nontrivial additional contributions to the conditional likelihood. In accordance with the previous spirit, we should decompose the path integral into contributions from the saddle points and their vicinity. We will now proceed to discuss these contributions individually.
@@ 4.2 The contributions from the saddle points
It is evident that both of these saddle-point equations are quadratic matrix equations, and therefore their exact solutions typically involve the application of some eigenvalue analysis techniques. Fortunately, in our case, the matrices $\mathcal{T}$ and $\mathcal{T}_{g}$ can be treated as perturbations, as they are proportional to the PNG parameter $f_{\rm NL}$. Next, we will discuss separately the computation of the matter likelihood and the joint likelihood, along with their leading-order contributions corresponding to PNG.
@@
By iteratively employing this method, we can obtain solutions to arbitrary higher orders. It is straightforward to observe that, after incorporating the leading-order PNG, the corrected saddle-point solution compared to eq. (4.13) effectively corresponds to a shift of the matter current $X\rightarrow X+\frac{iBf_{\rm NL}\delta^{2}}{D_{1}^{3}}$, and this correction is hierarchically suppressed by powers of $f_{\rm NL}$. This property enables us to carry over the discussion and results on the saddle-point contribution from [23], and to explore additional terms induced by the shift of $X$. Specifically, the terms in the action that give rise to corrections at order $f_{\rm NL}$ due to this adjustment are
@@
The aforementioned terms correspond to the supplementary contributions to the matter likelihood upon the introduction of PNG. Note that in calculating these contributions, we must consider not only the terms in the action that depend on $X$, but also the non-Gaussian effective interaction. By employing the same method, we can also extend this calculation to arbitrary higher orders.
@@
We now examine a more complicated scenario concerning the joint likelihood. Given the complexity of its formulation, we will first consider the simplest case, where $P_{\epsilon_{g}\epsilon_{m}}=0$, and subsequently restore the general case. By assuming the same field-decomposition ansatz and following an analogous derivation process, we can show that
$\bar{\boldsymbol{\phi}}_{g}=\boldsymbol{\phi}_{g,0}+\boldsymbol{\phi}_{g,1}\,,\;\boldsymbol{\phi}_{g,1}=-\frac{1}{2}(\mathcal{M}_{g}^{-1})^{ab}\mathcal{T}_{g}^{bcd}\boldsymbol{\phi}_{0,g}^{c}\boldsymbol{\phi}_{0,g}^{d}=\begin{pmatrix}0\\ \frac{iBf_{\rm NL}\delta^{2}}{2D_{1}^{3}}\\ 0\end{pmatrix}\,$. (4.20)
@@
By comparing eqs. (4.20) and (4.21), we can observe that after recovering $P_{\epsilon_{g}\epsilon_{m}}$, in addition to reproducing the same shift as before, all remaining terms are at least proportional to $P_{\epsilon_{g}\epsilon_{m}}$ or $f_{\rm NL}$. On the other hand, from our previous discussion, we know that $P_{\epsilon_{g}\epsilon_{m}}\sim k^{2}$, implying that its contribution is suppressed at large scales. This enables us to treat it as a perturbative parameter similar to $f_{\rm NL}$. Note that the leading-order saddle-point solution [24],
@@
inherently involves $P_{\epsilon_{g}\epsilon_{m}}$. Therefore, if we include the $P_{\epsilon_{g}\epsilon_{m}}$-related components in $\boldsymbol{\phi}_{g,1}$, many of the resulting terms naturally correspond to higher-order contributions. In the present work, our objective is to isolate the leading-order contributions proportional to $P_{\epsilon_{g}\epsilon_{m}}$ and $f_{\rm NL}$. Accordingly, it is reasonable to approximate the first-order perturbed solution as follows
@@
$X_{g,}$ $\rightarrow X_{g}-\frac{iBf_{\rm NL}P_{\epsilon_{g}\epsilon_{m}}\delta^{2}}{2P_{\epsilon_{g}}D_{1}^{3}}\,$, (4.25)
$\,X$ $\rightarrow X+\frac{Bf_{\rm NL}}{2D_{1}^{3}}\left(i\delta^{2}-2P_{\epsilon_{g}\epsilon_{m}}\delta X_{g,\rm cl}-\frac{ib_{1}P_{\epsilon_{g}\epsilon_{m}}\delta^{2}}{P_{\epsilon_{g}}}\right)\,$, (4.26)
@@
$\times\exp\left\{-\int_{\boldsymbol{k}}P_{\epsilon_{g}\epsilon_{m}}X_{g}X+\frac{f_{\rm NL}}{3!}\int_{\boldsymbol{k},\boldsymbol{p}_{1},\boldsymbol{p}_{2}}B(\delta_{\rm in})^{3}\right\}\bigg{|}_{\boldsymbol{\phi}_{g}=\bar{\boldsymbol{\phi}}_{g}}\,$, (4.27)
@@
A comparison of the matter and joint likelihood expressions shows that all irreducible saddle-point contributions are proportional to $f_{\rm NL}P_{\epsilon_{g}\epsilon_{m}}$, in agreement with the discussion in [23]. Because $P_{\epsilon_{g}\epsilon_{m}}\propto k^{2}$ corresponds to higher-derivative (gradient) terms and $f_{\rm NL}$ itself is small, these new terms are subject to a double suppression. Based on our previous estimates of the magnitudes of $f_{\rm NL}$ and $P_{\epsilon_{g}\epsilon_{m}}$, we find that the $P_{\epsilon_{g}\epsilon_{m}}$ contribution is typically subdominant to that of $f_{\rm NL}$. Consequently, the terms derived here are smaller than the $P_{\epsilon_{g}\epsilon_{m}}$ terms retained in [23] but larger than the neglected $P_{\epsilon_{g}\epsilon_{m}}^{2}$ contributions. Moreover, recent Bayesian forward-modeling analyses at the field level already include $f_{\rm NL}k^{2}$ terms – for example, the galaxy-bias expansion in [21] is carried out to this order. We therefore present these terms explicitly, both for completeness and as a pedagogical reference for future applications. At the same time, one might be concerned that certain terms in the action contain explicit factors of the imaginary unit $i$, potentially rendering the expression complex. However, this is not the case: due to the presence of $i$ in the definition of $X_{g,\rm cl}$ itself in eq. (4.22), all imaginary components in $S_{g}[\bar{\phi}_{g}]$ cancel out exactly, ensuring that the resulting expression is manifestly real – consistent with the physical interpretation of probability.
A brief comparison of the expressions for the two likelihoods with those under Gaussian initial conditions reveals certain underlying structures. We can observe that when $P_{\epsilon_{g}\epsilon_{m}}=0$, the contributions of PNG to both likelihoods are exactly the same, consistent with the discussion presented in [23]. Conversely, when the cross stochastic term is introduced, its contributions to the joint likelihood prevent this cancellation when computing the ratio of the two likelihoods, thereby imparting non-vanishing contributions to the conditional likelihood. The relationship between these noise terms and the likelihood expressions is of considerable interest, and we will discuss it in detail in the following subsection.
@@ 4.3 The contributions around the saddle points
We now turn to the section on precision calculations – specifically, employing the saddle-point expansion method to compute the contributions around the saddle points. To achieve this, we decompose the fields $\boldsymbol{\phi}$ and $\boldsymbol{\phi}_{g}$ into their saddle-point solutions and some small fluctuations around them
@@
It is evident that if the two eigenvalues are real, they must also be positive, as the growth factor $D_{1}$ can be consistently expected to be positive. Therefore, there are no concerns regarding negative-mode solutions at this stage. We can thus safely proceed with the saddle-point expansion method and apply the Gaussian integration formula, leading to the following result:
@@
Here, when plugging in $S[\bar{\boldsymbol{\phi}}]$, we have already incorporated the additional contributions arising from the inclusion of PNG. In this case, it is evident that although the eigenvalues $m_{n}$ exhibit explicit dependence on the matter field $\delta$, such dependence cancels out in their product. Consequently, the resulting matter likelihood precisely matches that obtained in [24]. As established in the previous discussions, this field-independent prefactor will be consistently absorbed into the normalization factor associated with the definition of functional Dirac delta function, and therefore it does not yield any nontrivial contribution to the matter likelihood.
@@
Even though we have employed such a redefinition to simplify our calculation, $S_{g}^{\prime\prime}[\bar{\boldsymbol{\phi}}_{g}]$ remains a three-dimensional matrix and may still possess negative or complex eigenvalues, thereby rendering the corresponding functional integral ill-defined. Accordingly, when computing the joint likelihood, it is essential to follow the strategy employed in section 2, complexifying the action and identifying the combination of the steepest descent contours in the complex plane that is homologous to the original integration contour (the real line in the field space). The full path integral can then be approximated by summing over all the contributions from these complex contours, which ensures a well-defined and convergent result. All relevant technical details have been provided in the preceding sections, and here we directly employ the gradient flow equation to determine the corresponding steepest descent contours. Assuming that the same steps as previously outlined have already been performed, the gradient flow equation in this case can be written as
@@
where we have used the fact that $\bar{\boldsymbol{\phi}}_{g}$ is independent of $u$, since $u$ simply serves as a parameter labeling different integration contours through the saddle point in the complex plane. At this time, the gradient flow equation is modified from its previous form eq. (2.7) by the addition of a non-Gaussian term, which is precisely equivalent to implementing the variable shift eq. (4.42) (note that $P_{\rm in}^{\prime\prime-1}$ is real). Therefore, by implementing this shift, one can construct the same eigen-equation and convert it into a Hermitian eigen-equation as in section 2, with the expressions remaining identical to those derived previously. Consequently, the sum over these positive and real eigenvalues $m_{n}$ can be obtained
@@
From this point forward, we will use the simplified notation $P_{\rm in}^{\prime\prime-1}$, with the understanding that it depends on the fields $\delta$ and $\delta_{g}$. With these elements established, we are now ready to compute the corresponding joint likelihood
@@
We can see that the resulting expression contains not only the saddle-point contributions but also an additional field-dependent prefactor. Furthermore, fluctuations around the saddle point generate a non-trivial contribution to the joint likelihood that is proportional to to $f_{\rm NL}$. This term is the leading-order correction sourced exclusively by PNG, and is therefore of particular importance. We will discuss it in detail in a later section. At this stage, the prefactor exhibits an explicit dependence on both the matter and galaxy fields, and cannot be eliminated by setting $P_{\epsilon_{g}\epsilon_{m}}=0$. In other words, once the contributions of PNG are taken into account, the precision calculation process introduces a field-dependent term that cannot be absorbed into the normalization procedure. Consequently, this term appears explicitly in the final expression for the conditional likelihood. Its potential implications will be examined in detail in the following subsection.